OpenStudy (anonymous):

will fan & medal :) Bruce has a bottle that contains 60% of lemon juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of lemon juice and water in the bottle. Define the variable used in the equation. Part B: How many liters of lemon juice are present in the bottle? Show your work.

OpenStudy (anonymous):

A: liters of lemon juice= (liters of water)/40*60 B: liters of lemon juice=1/40*60=1.5

OpenStudy (anonymous):

that explains it .. thank you no one would help me at all :) thank you again !